(1) A damped oscillator has a quality factor of 20. Part A:- By what fraction does the energy decrease during each cycle? Part B:- By what percentage does the damped angular frequency ωd differ from the undamped angular frequency?
Answer #1

Solution:
Part A

We know that fraction of energy decreasing per cycle is given by:

ΔE(t) / E(t) = 1 - e^(-2π/Q)

=>ΔE(t) / E(t) = 1 - e^(-2*3.14/20) = 1 - 0.73 = 0.27 Answer (A)